a man is 5 times as old as his son. after 2 years the man will be 4 times as old as his son find the present age of man
Answer by TimothyLamb(4379)

x = man age now
y = son age now
---
x = 5y
x + 2 = 4(y + 2)
---
put the system of linear equations into standard form
---
x = 5y
x + 2 = 4y + 8
---
x - 5y = 0
x - 4y = 6
---
copy and paste the above standard form linear equations in to this solver:
https://sooeet.com/math/system-of-linear-equations-solver.php
---
solution:
x = man age now = 30
y = son age now = 6
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations, quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Solve systems of linear equations up to 6-equations 6-variables:
https://sooeet.com/math/system-of-linear-equations-solver.php